# MoMaGen: Generating Demonstrations under Soft and Hard Constraints for Multi-Step Bimanual Mobile Manipulation

### v1

@@ 1 Introduction
Learning from human demonstrations is a powerful paradigm for teaching robots complex manipulation skills. A common approach for collecting such data is teleoperation, where a human directly controls the robot to demonstrate desired behaviors. When scaled up, teleoperated data has enabled training visuomotor policies with impressive generalization and success in challenging manipulation tasks (Brohan et al., 2022b; O’Neill et al., 2024; Khazatsky et al., 2024; Black et al., 2024; Intelligence et al., 2025). However, this data collection process remains expensive and time-consuming, especially for tasks that require high-quality demonstrations to ensure effective policy learning.
Recently, collecting a small amount of human teleoperation data and then synthesizing additional data in simulation has become a popular approach to scale up data collection (Mandlekar et al., 2023b; Garrett et al., 2024; Jiang et al., 2025b; Xue et al., 2025; Yang et al., 2025). Compared to offline data augmentation techniques such as those based on image augmentation (Laskin et al., 2020; Yarats et al., 2021; Pitis et al., 2020; Yu et al., 2023; Chen et al., 2023), this approach can autonomously generate new behaviorally diverse data for the same task. This process enlarges the support and convergence region of the policies and reduces the teacher-student distribution mismatch (Ross et al., 2011) through new generated experiences validated in simulation to ensure quality. Notably, the X-Gen family of techniques (Mandlekar et al., 2023b; Garrett et al., 2024; Jiang et al., 2025b; Xue et al., 2025; Yang et al., 2025) leverages simulation and augmentation based on a small number of human demonstrations that are used as seeds to generate multiple new variations automatically. While they have shown success in simple table-top manipulation tasks, a significant standing challenge for X-Gen methods is to extend the benefits of generalizable data generation to real-world tasks with more complex robot embodiments such as mobile manipulators.
@@
Solving real-world tasks, such as everyday household activities, often requires a mobile manipulator with whole-body control capabilities to coordinate stable and accurate navigation with end-effector manipulability, often with two arms (Jiang et al., 2025a; Li et al., 2020; Fu et al., 2024b). Teleoperation data collection becomes significantly challenging for high-degrees-of-freedom whole-body control since controlling the base and two arms is a severe overload on the human operators (Dass et al., 2024; Ishiguro et al., 2020; Fu et al., 2024b) (see Fig. 1, left). Augmenting a few (expensive) demonstrations becomes thus critical, but previous methods of the X-Gen family fall short in this domain due to two major reasons: First, mobile manipulation introduces the problem of object reachability. For novel object arrangements, naive replay of the navigation segments of the human collected demonstrations easily leads to robot configurations that render subsequent manipulation infeasible. Second, having a mobile base, and thus a movable camera, exacerbates the problems of partial observability. Concretely, when training visuomotor policies for mobile manipulation, a naive augmentation of demonstrations leads to severe problems in object visibility: the task-relevant objects may move out of the field of view, making it hard for the policy to make optimal decisions based on the images from onboard sensors. Naive motion planning (Garrett et al., 2024) or replay (Mandlekar et al., 2023a; Jiang et al., 2025b) is insufficient to ensure either reachability or visibility of the task-relevant objects.
@@ 2 Related Works
Data Acquisition for Robot Learning. Collecting large-scale human teleoperation data for robot learning incurs considerable costs. Scaling up data collection requires a large number of human operators over extended periods of time (Brohan et al., 2022b; O’Neill et al., 2024; Khazatsky et al., 2024). Offline data augmentation techniques can boost data quantity and diversity by perturbing existing trajectories (Mandlekar et al., 2023b; Garrett et al., 2024), or leveraging image augmentation techniques and generative models (Laskin et al., 2020; Yarats et al., 2021; Pitis et al., 2020; Yu et al., 2023; Chen et al., 2023). However, the augmented data may not always be executable by real robots. A promising alternative is to leverage automated data generation and validation in simulation. Fully automated approaches include trial-and-error (Levine et al., 2018; Pinto & Gupta, 2016; Yu et al., 2016; Dasari et al., 2019) and pre-programmed (e.g., scripted) experts (James et al., 2020; Gu et al., 2023; Jiang et al., 2022; Wang et al., 2023), which are yet to be proven effective for multi-step complex tasks with contact-rich interaction. X-Gen (Mandlekar et al., 2023b; Garrett et al., 2024; Jiang et al., 2025b; Xue et al., 2025; Yang et al., 2025) represents a hybrid approach that uses a handful of human demonstrations as seeds to generate many new variations, augmenting the data by a factor of $25\times$ to $350\times$ (Mandlekar et al., 2023b; Jiang et al., 2025b), while ensuring synthesized data is valid in simulation. A comparison between prior X-Gen works and our work can be found in Table 1. Our work signifies an important step toward a more generalizable data generation framework for challenging mobile manipulation tasks, which have never been tackled before.
Imitation Learning for Mobile Manipulation. Early successes in robot imitation learning mostly focused on fixed-based arms, but many real-world tasks require a mobile manipulator that can both navigate and manipulate. Such robots need to effectively navigate in an environment to position themselves for downstream manipulation. Collecting teleoperation data for mobile manipulation is significantly more costly: operators must simultaneously control the robot base and arms (Wong et al., 2022; Fu et al., 2024b; Shaw et al., 2024; Dass et al., 2024; Jiang et al., 2025a), calling for automated data generation methods for better scaling. On the algorithmic side, imitation learning methods started handle the complexities of mobile manipulation tasks, employing behavior cloning (Brohan et al., 2022a; Fu et al., 2024b; a; Cheng et al., 2024; Wu et al., 2024a; Yang et al., 2024; Li et al., 2024b; Ze et al., 2024; He et al., 2024) and large pretrained models (Ichter et al., 2022; Wu et al., 2023; Xu et al., 2023; Stone et al., 2023; Jiang et al., 2024b; Shah et al., 2024; Wu et al., 2024b) and imitating human videos (Bahety et al., 2025).
@@
MimicGen (Mandlekar et al., 2023b) Succ N/A
@@ 3 Problem Formulation: Automated Data Generation as Constrained Optimization
Following prior work (Mandlekar et al., 2023b; Garrett et al., 2024; Jiang et al., 2025b), each source demonstration $d$ can be decomposed to $N$ subtasks. Each subtask contains an object trajectory $S_{i}(o_{i}),i\in[N]$, where $o_{i}$ is the object of interest (i.e. target object), and an end-effector trajectory $\tau_{i}=\{\mathbf{T}^{E_{k}}_{W}\}_{k=0}^{K_{i}},i\in[N],k\in[K_{i}]$, where $\mathbf{T}^{E_{k}}_{W}$ is the pose of the end-effector frame $E$ with respect to the world reference frame $W$ at time $k$, and $K_{i}$ is the number of steps for the subtask $i$. Each subtask can further be labeled as either a free-space subtask or a contact-rich subtask. In a free-space subtask, the goal is to move the robot base or arms in free space (e.g. to a pregrasp pose), where feasible trajectories can be sampled using motion planning subject to kinematic and collision constraints. In a contact-rich subtask, the goal is to manipulate the relevant objects through contacts (e.g. grasping, wiping). The relative poses between the end-effector frame and the target object frame for contact-rich subtasks are preserved from the source to the generated demonstrations.
@@ 4.1 Constraints for Bimanual Mobile Manipulation
Reachability as Hard Constraint. A key distinction of mobile manipulation is the need to control the robot base for effective downstream manipulation. Prior works (Mandlekar et al., 2023a; Jiang et al., 2025b) reuse base trajectories from source demos without adaptation, which fails when randomized target objects lie outside the arm’s workspace. To address this, we impose reachability as a hard constraint, ensuring sampled base poses keep all required end-effector trajectories within reach.
@@ 5.1 Task Setup
We evaluate MoMaGen on four household tasks that require mobile manipulation and involve various manipulation skills, including pick-and-place, contact-rich interactions and bimanual coordination, illustrated in Figure 3. Inspired by BEHAVIOR-1K, all tasks are implemented in OmniGibson (Li et al., 2023; 2024a). Pick Cup: navigate to a table and lift a cup. Tidy Table: move a cup from the countertop to the sink, requiring long-range mobile manipulation. Put Dishes Away: stack two plates from the countertop onto a shelf using two arms independently, requiring bimanual uncoordinated motion. Clean Frying Pan: use both arms to scrub a dusty pan with a brush, requiring contact-rich bimanual coordinated motion. For each task, we collect a single 1–3 minute human teleoperated demonstration, with base motion accounting for 45% of the duration.
@@ 5.2 Domain Randomization Schemes
For each task, we define three domain randomization levels of increasing difficulty. D0: task-relevant objects randomized within $\pm 15$ cm and $\pm 15^{\circ}$ on the same furniture. D1: task-relevant objects placed anywhere on the furniture with unrestricted orientation (Figure 4a). D2: D1 plus additional objects on furniture (manipulation obstacles) and on the floor (navigation obstacles) (Figure 1, upper right). This scheme is significantly more aggressive than prior work (Mandlekar et al., 2023b; Garrett et al., 2024; Jiang et al., 2025b), enabled by MoMaGen ’s unique ability to generate novel base motions. More details and visualizations can be found in Sec. B.3 and Fig. 9 in the Appendix.
@@ 5.3 Data Generation Comparison
We evaluate MoMaGen for bimanual mobile manipulation under varying randomization schemes, comparing it to two baselines: SkillMimicGen (Garrett et al., 2024), which generates single-arm trajectories via motion planning and task-space control, and DexMimicGen (Jiang et al., 2025b), which targets dexterous bimanual data. Since all the evaluated tasks require substantial base movement, both baselines are extended with base trajectory replay from source demos, similar to MimicGen (Mandlekar et al., 2023a). We assess methods using three metrics: (1) data diversity (object pose and action variation), (2) generation success rate, and (3) object visibility ratio during navigation.
@@
How diverse are the demonstrations generated by MoMaGen? Figure 4a shows the variation in task-relevant object poses for the Tidy Table task under MoMaGen (D0/D1) and SkillMimicGen (D0). The baselines succeed only in D0, as they cannot generate novel base motions. In contrast, MoMaGen (D1) covers a much wider range of object poses, spanning the entire counter rather than a small corner. Figures 4b–c further demonstrate that MoMaGen D1 yields substantially broader coverage of base and end-effector poses compared to D0 and the baselines. Figure 4d shows PCA projections of arm and torso joints under D0. Despite identical object pose distributions, randomizing feasible base placements allows MoMaGen to achieve far greater action diversity. Moreover, with its motion planner, MoMaGen can generate demonstrations in cluttered scenes with distractors or obstacles (Figure 9), further enhancing the diversity of both robot actions and visual observations. Additional results for other tasks are shown in Figures 10, 11, and 12 in the Appendix.
@@
Can MoMaGen generate demos with a new robot embodiment? We evaluate cross-embodiment data generation by using a single source demo from a Galexea R1 robot to generate Pick Cup demos on a TIAGo robot. MoMaGen successfully transfers across platforms by planning/replaying dense end-effector trajectories in the task space, which is largely agnostic to robot-specific kinematic, demonstrating robustness and flexibility of MoMaGen. More discussions are found in Appendix B.2.
@@ 5.4 Policy Learning with Generated Demonstrations
Figure 7: Data Scaling of $\pi_{0}$.
@@ 6 Conclusions and Limitations
We provide all the source code for simulation task setups, data generation, policy training, and detailed instructions for reproducing all data generation and policy learning results in our code submission (included in the supplementary materials). All training hyperparameters and implementation details are documented in Appendix C to facilitate full reproducibility.
@@ A.1 Sim-to-real
Experiment setup. Zero-shot transfer is challenging due to domain gaps, particularly for vision-based policies (Jiang et al., 2024a). In our case, simulated and real RGB images differ substantially, and real-world depth estimates are noisy. To mitigate this, we collect 50 real-world demonstrations for the Pick Cup (D0) task and evaluate whether simulation pretraining improves real-world learning.

### v2

@@ 1 Introduction
Learning from human demonstrations is a powerful paradigm for teaching robots complex manipulation skills. A common approach for collecting such data is teleoperation, where a human directly controls the robot to demonstrate desired behaviors. When scaled up, teleoperated data has enabled training visuomotor policies with impressive generalization and success in challenging manipulation tasks (Brohan et al., 2022; O’Neill et al., 2024; Khazatsky et al., 2024; Black et al., 2024; Intelligence et al., 2025). However, this data collection process remains expensive and time-consuming, especially for tasks that require high-quality demonstrations to ensure effective policy learning.
Recently, collecting a small amount of human teleoperation data and then synthesizing additional data in simulation has become a popular approach to scale up data collection (Mandlekar et al., 2023; Garrett et al., 2024; Jiang et al., 2025b; Xue et al., 2025; Yang et al., 2025). Compared to offline data augmentation techniques such as those based on image augmentation (Laskin et al., 2020; Yarats et al., 2021; Pitis et al., 2020; Yu et al., 2023; Chen et al., 2023), this approach can autonomously generate new behaviorally diverse data for the same task. This process enlarges the support and convergence region of the policies and reduces the teacher-student distribution mismatch (Ross et al., 2011) through new generated experiences validated in simulation to ensure quality. Notably, the X-Gen family of techniques (Mandlekar et al., 2023; Garrett et al., 2024; Jiang et al., 2025b; Xue et al., 2025; Yang et al., 2025) leverages simulation and augmentation based on a small number of human demonstrations that are used as seeds to generate multiple new variations automatically. While they have shown success in simple table-top manipulation tasks, a significant standing challenge for X-Gen methods is to extend the benefits of generalizable data generation to real-world tasks with more complex robot embodiments such as mobile manipulators.
@@
Solving real-world tasks, such as everyday household activities, often requires a mobile manipulator with whole-body control capabilities to coordinate stable and accurate navigation with end-effector manipulability, often with two arms (Jiang et al., 2025a; Li et al., 2020; Fu et al., 2024). Teleoperation data collection becomes significantly challenging for high-degrees-of-freedom whole-body control since controlling the base and two arms is a severe overload on the human operators (Dass et al., 2024; 9158331; Fu et al., 2024) (see Fig. 1, left). Augmenting a few (expensive) demonstrations becomes thus critical, but previous methods of the X-Gen family fall short in this domain due to two major reasons: First, mobile manipulation introduces the problem of object reachability. For novel object arrangements, naive replay of the navigation segments of the human collected demonstrations easily leads to robot configurations that render subsequent manipulation infeasible. Second, having a mobile base, and thus a movable camera, exacerbates the problems of partial observability. Concretely, when training visuomotor policies for mobile manipulation, a naive augmentation of demonstrations leads to severe problems in object visibility: the task-relevant objects may move out of the field of view, making it hard for the policy to make optimal decisions based on the images from onboard sensors. Naive motion planning (Garrett et al., 2024) or replay (mandlekar2023mimicgen; Jiang et al., 2025b) is insufficient to ensure either reachability or visibility of the task-relevant objects.
@@ 2 Related Works
Data Acquisition for Robot Learning. Collecting large-scale human teleoperation data for robot learning incurs considerable costs. Scaling up data collection requires a large number of human operators over extended periods of time (Brohan et al., 2022; O’Neill et al., 2024; Khazatsky et al., 2024). Offline data augmentation techniques can boost data quantity and diversity by perturbing existing trajectories (Mandlekar et al., 2023; Garrett et al., 2024), or leveraging image augmentation techniques and generative models (Laskin et al., 2020; Yarats et al., 2021; Pitis et al., 2020; Yu et al., 2023; Chen et al., 2023). However, the augmented data may not always be executable by real robots. A promising alternative is to leverage automated data generation and validation in simulation. Fully automated approaches include trial-and-error (Levine et al., 2018; Pinto and Gupta, 2016; Yu et al., 2016; Dasari et al., 2019) and pre-programmed (e.g., scripted) experts (James et al., 2020; Gu et al., 2023; Jiang et al., 2022; Wang et al., 2023), which are yet to be proven effective for multi-step complex tasks with contact-rich interaction. X-Gen (Mandlekar et al., 2023; Garrett et al., 2024; Jiang et al., 2025b; Xue et al., 2025; Yang et al., 2025) represents a hybrid approach that uses a handful of human demonstrations as seeds to generate many new variations, augmenting the data by a factor of $25\times$ to $350\times$ (Mandlekar et al., 2023; Jiang et al., 2025b), while ensuring synthesized data is valid in simulation. A comparison between prior X-Gen works and our work can be found in Table 1. Our work signifies an important step toward a more generalizable data generation framework for challenging mobile manipulation tasks, which have never been tackled before.
Imitation Learning for Mobile Manipulation. Early successes in robot imitation learning mostly focused on fixed-based arms, but many real-world tasks require a mobile manipulator that can both navigate and manipulate. Such robots need to effectively navigate in an environment to position themselves for downstream manipulation. Collecting teleoperation data for mobile manipulation is significantly more costly: operators must simultaneously control the robot base and arms (Wong et al., 2022; Fu et al., 2024; Shaw et al., 2024; Dass et al., 2024; Jiang et al., 2025a), calling for automated data generation methods for better scaling. On the algorithmic side, imitation learning methods started handle the complexities of mobile manipulation tasks, employing behavior cloning (brohan2022rt1; Fu et al., 2024; fu2024humanplus; cheng2024opentelevision; wu2024tidybot; yang2024equibot; li2024okami; ze2024humanoid_manipulation; he2024omniho) and large pretrained models (DBLP:conf/corl/IchterBCFHHHIIJ22; wu23tidybot; xu2023creative; stone2023openworld; jiang2024harmon; shah2024bumble; wu2024helpful).
@@
MimicGen (Mandlekar et al., 2023) Succ N/A
@@ 3 Problem Formulation: Automated Data Generation as Constrained Optimization
Following prior work (Mandlekar et al., 2023; Garrett et al., 2024; Jiang et al., 2025b), each source demonstration $d$ can be decomposed to $N$ subtasks. Each subtask contains an object trajectory $S_{i}(o_{i}),i\in[N]$, where $o_{i}$ is the object of interest (i.e. target object), and an end-effector trajectory $\tau_{i}=\{\mathbf{T}^{E_{k}}_{W}\}_{k=0}^{K_{i}},i\in[N],k\in[K_{i}]$, where $\mathbf{T}^{E_{k}}_{W}$ is the pose of the end-effector frame $E$ with respect to the world reference frame $W$ at time $k$, and $K_{i}$ is the number of steps for the subtask $i$. Each subtask can further be labeled as either a free-space subtask or a contact-rich subtask. In a free-space subtask, the goal is to move the robot base or arms in free space (e.g. to a pregrasp pose), where feasible trajectories can be sampled using motion planning subject to kinematic and collision constraints. In a contact-rich subtask, the goal is to manipulate the relevant objects through contacts (e.g. grasping, wiping). The relative poses between the end-effector frame and the target object frame for contact-rich subtasks are preserved from the source to the generated demonstrations.
@@ 4.1 Constraints for Bimanual Mobile Manipulation
Reachability as Hard Constraint. A key distinction of mobile manipulation is the need to control the robot base for effective downstream manipulation. Prior works (mandlekar2023mimicgen; Jiang et al., 2025b) reuse base trajectories from source demos without adaptation, which fails when randomized target objects lie outside the arm’s workspace. To address this, we impose reachability as a hard constraint, ensuring sampled base poses keep all required end-effector trajectories within reach.
@@ 5.1 Task Setup
We evaluate MoMaGen on four household tasks that require mobile manipulation and involve various manipulation skills, including pick-and-place, contact-rich interactions and bimanual coordination, illustrated in Figure 3. Inspired by BEHAVIOR-1K, all tasks are implemented in OmniGibson (Li et al., 2023; 2024). Pick Cup: navigate to a table and lift a cup. Tidy Table: move a cup from the countertop to the sink, requiring long-range mobile manipulation. Put Dishes Away: stack two plates from the countertop onto a shelf using two arms independently, requiring bimanual uncoordinated motion. Clean Frying Pan: use both arms to scrub a dusty pan with a brush, requiring contact-rich bimanual coordinated motion. For each task, we collect a single 1–3 minute human teleoperated demonstration, with base motion accounting for 45% of the duration.
@@ 5.2 Domain Randomization Schemes
For each task, we define three domain randomization levels of increasing difficulty. D0: task-relevant objects randomized within $\pm 15$ cm and $\pm 15^{\circ}$ on the same furniture. D1: task-relevant objects placed anywhere on the furniture with unrestricted orientation (Figure 4a). D2: D1 plus additional objects on furniture (manipulation obstacles) and on the floor (navigation obstacles) (Figure 1, upper right). This scheme is significantly more aggressive than prior work (Mandlekar et al., 2023; Garrett et al., 2024; Jiang et al., 2025b), enabled by MoMaGen ’s unique ability to generate novel base motions. More details and visualizations can be found in Sec. B.3 and Fig. 12 in the Appendix.
@@ 5.3 Data Generation Comparison
We evaluate MoMaGen for bimanual mobile manipulation under varying randomization schemes, comparing it to two baselines: SkillMimicGen (Garrett et al., 2024), which generates single-arm trajectories via motion planning and task-space control, and DexMimicGen (Jiang et al., 2025b), which targets dexterous bimanual data. Since all the evaluated tasks require substantial base movement, both baselines are extended with base trajectory replay from source demos, similar to MimicGen (mandlekar2023mimicgen). We assess methods using three metrics: (1) data diversity (object pose and action variation), (2) generation success rate, and (3) object visibility ratio during navigation.
@@
How diverse are the demonstrations generated by MoMaGen? Figure 4a shows the variation in task-relevant object poses for the Tidy Table task under MoMaGen (D0/D1) and SkillMimicGen (D0). The baselines succeed only in D0, as they cannot generate novel base motions. In contrast, MoMaGen (D1) covers a much wider range of object poses, spanning the entire counter rather than a small corner. Figures 4b–c further demonstrate that MoMaGen D1 yields substantially broader coverage of base and end-effector poses compared to D0 and the baselines. Figure 4d shows PCA projections of arm and torso joints under D0. Despite identical object pose distributions, randomizing feasible base placements allows MoMaGen to achieve far greater action diversity. Moreover, with its motion planner, MoMaGen can generate demonstrations in cluttered scenes with distractors or obstacles (Figure 12), further enhancing the diversity of both robot actions and visual observations. Additional results for other tasks are shown in Figures 13, 14, and 15 in the Appendix.
@@
Can MoMaGen generate demos with a new robot embodiment? We evaluate cross-embodiment data generation by using a single source demo from a Galexea R1 robot to generate Pick Cup demos on a TIAGo robot. MoMaGen successfully transfers across platforms by planning/replaying dense end-effector trajectories in the task space, which is largely agnostic to robot-specific kinematics, demonstrating robustness and flexibility of MoMaGen. More discussions are found in Appendix B.2.
@@ 5.4 Policy Learning with Generated Demonstrations
Figure 7: Data Scaling of $\pi_{0}$
@@
5.5 MoMaGen  Failure and Computational Cost Analysis
Figure 8: Failure Analysis
Failure Analysis. We analyze the causes of unsuccessful data generation episodes by categorizing failures into base sampling, base-level planning, arm-level planning, and simulation instabilities. Base Sampling failures occur when no base pose is found that satisfies reachability and visibility constraints. Base MP IK and Base MP TrajOpt correspond to inverse-kinematics and trajectory-optimization failures during base motion planning. Similarly, Arm MP IK and Arm MP TrajOpt capture IK and trajectory-optimization failures during arm motion planning, while Arm MP Other includes additional arm-planning issues such as invalid start or goal states. Simulation Instabilities arise from controller inaccuracies and stochastic effects in the simulator. Fig. 8 shows that across all three domain randomizations, simulation instabilities account for a substantial share of failures (35%), which can be mitigated by improvements in controller robustness and simulation fidelity. Arm-level motion planning contributes the largest proportion of planner-related failures (40% on average), exceeding base-level planning failures (26% on average). Notably, in D2 randomization, navigation-related failures (base sampling, base IK, and base TrajOpt) increase significantly due to floor obstacles in an already tight navigation space. We also analyze the distribution of failures across task-level steps within each multi-step task (Fig. 16 in Appendix).
Figure 9: Compute costs (wall-clock time) for MoMaGen data generation components.
Figure 10: Compute cost of each component as a percentage of the episode
Compute Cost Analysis. To better understand the efficiency of our data generation pipeline, we analyze the computational cost contribution of each component. Fig. 9 shows the wall-clock time distributions for base sampling, motion planning, and simulation execution. Simulation execution dominates total computation time, greatly exceeding the corresponding planning durations. For instance, base motion planning averages 18 seconds, whereas executing the resulting motion in simulation takes 100 seconds. Base sampling also exhibits high variance due to our current strategy of random sampling within a ring-shaped region around the target: when only a few poses are feasible, the search becomes slow, whereas abundant feasible poses lead to quick success. Future work may reduce this variance through more informed sampling methods—for example, biasing toward regions with greater free space. Fig. 10 reports the average compute cost of each component as a percentage of the total episode duration, computed only over successful episodes to ensure a fair comparison across all components. Notably, base sampling becomes increasingly expensive from D0 to D2—reflecting the rise in scene complexity, where fewer feasible base poses exist and longer searches are required.
@@ 6 Conclusions and Limitations
We provide all the source code for simulation task setups, data generation, policy training, and detailed instructions for reproducing all data generation and policy learning results in our code submission (included in the supplementary materials) as well as on our website. All training hyperparameters and implementation details are documented in Appendix C to facilitate full reproducibility.
@@ A.1 Sim-to-real
Experiment setup. Zero-shot transfer is challenging due to domain gaps, particularly for vision-based policies (jiang2024transic). In our case, simulated and real RGB images differ substantially, and real-world depth estimates are noisy. To mitigate this, we collect 50 real-world demonstrations for the Pick Cup (D0) task and evaluate whether simulation pretraining improves real-world learning.